### v2

@@ Abstract
Colonoscopic Polyp Re-Identification aims to match the same polyp from a large gallery with images from different views taken using different cameras, which plays an important role in the prevention and treatment of colorectal cancer in computer-aided diagnosis. However, traditional methods for object ReID directly adopting CNN models trained on the ImageNet dataset usually produce unsatisfactory retrieval performance on colonoscopic datasets due to the large domain gap. Worsely, these solutions typically learn unimodal modal representations on the basis of visual samples, which fails to explore complementary information from other different modalities. To address this challenge, we propose a novel Deep Multimodal Collaborative Learning framework named DMCL for polyp re-identification, which can effectively encourage modality collaboration and reinforce generalization capability in medical scenarios. On the basis of it, a dynamic multimodal feature fusion strategy is introduced to leverage the optimized multimodal representations for multimodal fusion via end-to-end training. Experiments on the standard benchmarks show the benefits of the multimodal setting over state-of-the-art unimodal ReID models, especially when combined with the specialized multimodal fusion strategy, from which we have proved that learning representation with multiple-modality can be competitive to methods based on unimodal representation learning. We also hope that our method will shed light on some related researches to move forward, especially for multimodal collaborative learning. The code is publicly available at https://github.com/JeremyXSC/DMCL.
@@
Colonoscopic Polyp Re-Identification, modal representation, multimodal collaborative learning, generalization capability
@@ 1 Introduction
Colonoscopic polyp re-identification (Polyp ReID) aims to match a specific polyp in a large gallery with different cameras and locations, which has been studied intensively due to its practical importance in the prevention and treatment of colorectal cancer in the computer-aided diagnosis (Koffas et al, 2022). With the development of deep convolution neural networks and the availability of video re-identification dataset, video retrieval methods have achieved remarkable performance in a supervised manner (Feng et al, 2019; Xiang et al, 2023b, c), where a model is trained and tested on different splits of the same dataset. However, in practice, manually labeling a large diversity of pairwise polyp area data is time-consuming and labor-intensive when directly deploying polyp ReID system to new hospital scenarios (Chen et al, 2023). Nevertheless, when compared with the conventional ReID (Xiang et al, 2020c, a), polyp ReID is confronted with more challenges in some aspects: 1) from the model perspective: traditional object ReID methods learn the unimodal representation by greedily “pre-training” several layers of features on the basis of visual samples (Xiang et al, 2023a, 2024a), while ignore to explore complementary information from different modalities, and 2) from the data perspective, polyp ReID will encounter many challenges such as variation in terms of backgrounds, viewpoint, and illumination, etc., which poses great challenges to the clinical deployment of deep model in real-world scenarios.
Based on the aforementioned findings, we novely propose a deep multimodal collaborative learning framework named DMCL to encourage modality collaboration and reinforce generalization capability in medical scenarios. On the basis of it, a dynamic multimodal training strategy is introduced to leverage the unimodal representations for multimodal fusion via end-to-end training on multimodal tasks, and then improve the performance of our model in an end-to-end manner, shown in Fig. 1. To the best of our knowledge, this is the first attempt to employ the visual-text feature with collaborative training mechanism for colonoscopic polyp re-identification.
@@
    A deep multimodal collaborative learning framework DMCL is proposed for the first time to obtain the visual and texture information simultaneously, and then promote the development of multimodal polyp re-identification.
@@
    Based on the DMCL framework, we introduce a simple but effective multimodal feature fusion strategy, which can greatly help the model learn more discriminatve information during training and texting stage.
@@
    Comprehensive experiments on standard benchmark datasets demonstrate that our method can achieve state-of-the-art results, proving the efficacy of our method.
@@ 2 Related Work
In this section, we have a brief review on the related works of traditional re-identification methods. The mainstream idea of the existing methods is to learn a discriminative and robust model for downstream polyp ReID task. There methods can be roughly divided into hand-crafted based approaches and deep learning based approaches.
@@ 2.1 Hand-crafted based Approaches
Traditional research works (Prosser et al, 2010; Zhao et al, 2013, 2014) related to hand-crafted systems for image retrieval task aim to design or learn discriminative representation or pedestrian features. For example, Prosser et al (2010) propose a reformulation of the person re-identification problem as a learning to rank problem. Zhao et al (2013) exploit the pairwise salience distribution relationship between pedestrian images, and solve the person re-identification problem by proposing a salience matching strategy. Besides directly using mid-level color and texture features, some methods (Zhao et al, 2014; Xiang et al, 2020a) also explore different discriminative abilities of local patches for robust generalization. Unfortunately, these handcrafted feature based approaches always fail to produce competitive results on large-scale datasets. The main reason is that these early works are mostly based on heuristic design, and thus they could not learn optimal discriminative features on current large-scale datasets.
@@ 2.2 Deep Learning based Approaches
Recently, there has been a significant research interest in the design of deep learning based approaches for image or video retrieval (Shao et al, 2021; Xiang et al, 2023a; Ma et al, 2020; Xiang et al, 2020a). For example, Shao et al (2021) propose temporal context aggregation which incorporates long-range temporal information for content-based video retrieval. Xiang et al (2020a) propose a feature fusion strategy based on traditional convolutional neural network for pedestrian retrieval. Ma et al (2020) explore a method for performing video moment retrieval in a weakly-supervised manner. Lin et al (2020) propose a semantic completion network including the proposal generation module to score all candidate proposals in a single pass. Besides, Kordopatis-Zilos et al (2022) propose a video retrieval framework based on knowledge distillation that addresses the problem of performance-efficiency trade-off. As for the self-supervised learning, Wu et al (2018) present an unsupervised feature learning approach called instance-wise contrastive learning by maximizing distinction between instances via a non-parametric softmax formulation. In addition, Chen et al (2023) propose a self-supervised contrastive representation learning scheme named Colo-SCRL to learn spatial representations from colonoscopic video dataset. However, on the one hand, all above approaches divide the learning process into multiple stages, each requiring independent optimization. On the other hand, despite the tremendous successes achieved by deep learning-based approach, they have been left largely unexplored in terms of semantic feature embedding.
In this work, we propose a novel Deep Multimodal Collaborative Learning framework named DMCL for multimodal polyp re-identification, which remarkably encourage modality collaboration and reinforce generalization capability in medical scenarios. On the basis of it, a dynamic multimodal feature fusion strategy is proposed to fully exploit the prior knowledge of visual-text feature, thereby improving the accuracy of polyp ReID in complex textured regions of unencoded subimages. To the best of our knowledge, this is the first attempt to employ the visual-text feature with collaborative learning mechanism for colonoscopic polyp re-identification task. To this end, the deep multimodal collaborative learning system proposed in this study not only improves the polyp ReID performance to a certain extent but also effectively maintains the social logical consistency of the extracted feature, which ensures the reliability and usability of the paradigm in various application scenarios, providing strong support for its application.
@@ 3.1 Preliminary
Figure 2: The overview of our proposed method on polyp re-identification task, which contains two main parts: 1) the visual feature backbone and 2) texture feature backbone respectively. Specially, visual feature backbone is composed of CNN network, and texture feature backbone is consisted of Transformer architecture named Bert model.
@@ 3.2 Our Proposed DMCL network
In this work, we design a DMCL framework for polyp re-identification in medical scenarios, which includes a visual feature extractor, a textual feature encoder and dynamic multimodal feature fusion strategy.
@@
Text Encoder. As for the text encoder module, we employ the BERT model (Devlin et al, 2018) to encode the text description of corresponding polyp. In essence, BERT model is a pre-trained language model based on the Transformer architecture, which learns language representations through masked language modeling and next sentence prediction tasks on a large amount of unlabeled text. In this work, given a text description of the polyp with $n$ characters, BERT will encode it into a $n$$\times$ 768-dimensions vector matrix for subsequent feature fusion and processing, formulated as $[T_{1},T_{2},T_{3},\cdots,T_{n}]^{1\times 768}$. The bidirectional encoder design of BERT enables it to capture contextual information of words, generating more accurate text representations.
Dynamic Multimodal Feature Fusion. In the general scenarios, there exists the obvious data discrepancy between image and text in our dataset, which inevitably has a negative impacts on the model’s performance. To overcome this issue, we creatively introduce a multimodal fusion strategy based on the self-attention mechanism, which is a crucial part of our proposed deep collaborative training framework. To be more specific, we obtain the image encoding of $I$ after passing through the image encoder, whose size is then dimensionally reduced from the original 2048-dimension to 768-dimension. Similarly, we can also obtain the text encoding through the text encoder, which is denoted as:
@@
$K=[I,T_{1},T_{2},T_{3},\cdots,T_{n}]$ (2)
where $K$ represents the features generated by the multimodal fusion component. Remarkably, multimodal fusion component consists of multiple self-attention layers to accomplish the fusion of visual feature and text feature. After passing through three linear layers, these three vectors are then inputted into the self-attention module. The output from the self-attention module is:
[…]
where $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ represents the Query, Key, and Value of self-attention mechanism respectively, and $d$ denotes the dimension of the input vector. After modal fusion, we can obtain the new updated features:
$K_{\text{f}}=\left[I_{f0},T_{f1},T_{f2},T_{f3},\ldots,T_{fn}\right]$ (4)
Generally speaking, visual feature is of paramount importance for the task of polyp re-identification. Inspired by this, we choose the feature of $I_{f0}$ at the corresponding position of the original image feature, which can be adopted as the testing feature of the corresponding polyp in the testing phase for downstream polyp ReID task. The whole procedure of our collaborative multimodal training system is illustrated in Algorithm 1.
@@
12:     Perform multimodal feature fusion: $K_{\text{f}}\leftarrow I+T$;
@@
15:Performing testing with model $\theta_{opt}$;
@@ 3.3 Dynamic Network Updating
In essence, many previous works (Luo et al, 2019; Xiang et al, 2023a; Luo et al, 2019; Xiang et al, 2020b) have been found that performing training with multiple losses has great potential to learn a robust and generalizable ReID model. During the training process, we adopt a identification loss (Zheng et al, 2017a) and triplet loss (Hermans et al, 2017) as the optimization metric for our DMCL model. Specially, the triplet-loss function aims to reduce the feature distance between similar polyp images while increasing the feature distance between different polyp images. The identification loss function, on the other hand, is to enhance the model’s ability to recognize and classify polyp categories.
To be more specific, for a single-label $N$ classification task, the identification loss (cross-entropy loss) is written as,
$\mathcal{L}_{ID}=-\frac{1}{\text{ $M_{batch}$ }}\sum_{i=1}^{\text{ $M_{batch}$}}\sum_{j=1}^{N}y_{ij}\log\hat{y_{ij}}$ (5)
@@ 4.1 Datasets and Evaluation Metric
We conduct experiments on several large-scale public datasets, which include Colo-Pair (Chen et al, 2023), Market-1501 (Zheng et al, 2015), DukeMTMC-reID (Zheng et al, 2017b) and CUHK03 dataset (Li et al, 2014). We follow the standard evaluation protocol (Zheng et al, 2015) used in the ReID task and adopt mean Average Precision (mAP) and Cumulative Matching Characteristics (CMC) at Rank-1, Rank-5, and Rank-10 for performance evaluation on downstream ReID task.
@@
ViSiL (Kordopatis-Zilos et al, 2019) ICCV 19 24.9 14.5 30.6 51.6
CoCLR (Han et al, 2020) NIPS 20 16.3 6.5 22.6 33.9
TCA (Shao et al, 2021) WCAV 21 27.8 16.1 35.5 53.2
ViT (Caron et al, 2021) CVPR 21 20.4 9.7 30.6 43.5
CVRL (Qian et al, 2021) CVPR 21 23.6 11.3 32.3 53.2
CgSc (Kordopatis-Zilos et al, 2022) IJCV 22 21.4 8.1 35.5 45.2
@@
Colo-SCRL (Chen et al, 2023) ICME 23 31.5 22.6 41.9 58.1
VT-ReID (Xiang et al, 2024b) ICASSP 24 37.9 23.4 44.5 60.1
@@
4.2 Implementation details
In our experiment, ResNet-50 (He et al, 2016) is regarded as the backbone with no bells and whistles during visual feature extraction. Following the training procedure in (Tong et al, 2022), we adopt the common methods such as random flipping and random cropping for data augmentation and employ the Adam optimizer with a weight decay co-efficient of $1\times 10^{-5}$ and $1\times 10^{-7}$ for parameter optimization. Besides, we adopt the ID loss and triplet loss functions to train the model for 180 iterations, and the cosine distance is also adopted to calculate the similarity of polyp features in the dataset for the task of polyp re-identification. Please note that the text information is also employed during testing phase. In addition, the batch size $M_{batch}$ for training is set to 64. All the experiments are performed on PyTorch framework with one Nvidia GeForce RTX 2080Ti GPU on a server equipped with an Intel Xeon Gold 6130T CPU.

### v3

@@ Abstract
Colonoscopic Polyp Re-Identification aims to match the same polyp from a large gallery with images from different views taken using different cameras, which plays an important role in the prevention and treatment of colorectal cancer in computer-aided diagnosis. However, traditional methods for object ReID directly adopting CNN models trained on the ImageNet dataset usually produce unsatisfactory retrieval performance on colonoscopic datasets due to the large domain gap. Worsely, these solutions typically learn unimodal modal representations on the basis of visual samples, which fails to explore complementary information from other different modalities. To address this challenge, we propose a novel Deep Multimodal Collaborative Learning framework named DMCL for polyp re-identification, which can effectively encourage multimodal knowledge collaboration and reinforce generalization capability in medical scenarios. On the basis of it, a dynamic multimodal feature fusion strategy is introduced to leverage the optimized visual-text representations for multimodal fusion via end-to-end training. Experiments on the standard benchmarks show the benefits of the multimodal setting over state-of-the-art unimodal ReID models, especially when combined with the collaborative multimodal fusion strategy. The code is publicly available at https://github.com/JeremyXSC/DMCL.
@@
colonoscopic polyp re-identification, multimodal collaborative learning, generalization capability
@@ 1 Introduction
Colonoscopic polyp re-identification (Polyp ReID) aims to match a specific polyp in a large gallery with different cameras and locations, which has been studied intensively due to its practical importance in the prevention and treatment of colorectal cancer in the computer-aided diagnosis (Chen et al, 2023), it also provides critical support for clinical applications including longitudinal lesion monitoring, diagnostic decision-making, and therapeutic outcome assessment. Specially, the task diagram of colonoscopic polyp re-identification is illustrated in Fig. 1. With the development of deep convolution neural networks and the availability of video re-identification dataset, video retrieval methods have achieved remarkable performance in a supervised manner (Feng et al, 2019; Xiang et al, 2023b), where a model is trained and tested on different splits of the same dataset. However, in practice, manually labeling a large diversity of pairwise polyp area data is time-consuming and labor-intensive when directly deploying polyp ReID system to new hospital scenarios. Nevertheless, when compared with the conventional ReID, polyp ReID is confronted with more challenges in some aspects: 1) from the model perspective: traditional object ReID methods learn the unimodal representation by greedily “pre-training” several layers of features on the basis of visual samples, while ignore to explore complementary information from different modalities, and 2) from the data perspective, polyp ReID will encounter many challenges such as variation in terms of backgrounds, viewpoint, and illumination, etc., which poses great challenges to the clinical deployment of deep model in real-world scenarios (Xiang et al, 2025).
Figure 1: The task diagram of colonoscopic polyp re-identification, which matches and correlates polyps appearing across different temporal frames, viewing angles, or clinical examinations. The core technical challenge involves determining whether polyps visualized in distinct images or video sequences represent the same pathological lesion.
Based on the aforementioned findings, we novelly propose a Deep Multimodal Collaborative Learning framework named DMCL to encourage multimodal knowledge collaboration and reinforce generalization capability in medical scenarios. On the basis of it, a dynamic multimodal training strategy is introduced to leverage the unimodal representations for multimodal fusion via end-to-end training on multimodal tasks, and then improve the performance of our model in an end-to-end manner. The illustration of our multimodal polyp dataset is shown in Fig. 2. To the best of our knowledge, this is the first attempt to employ the visual-text feature with collaborative training mechanism for colonoscopic polyp re-identification, from which we have proved that learning representation with multiple-modality can be competitive to methods based on unimodal representation learning. We also hope that our method will shed light on some related research to move forward, especially for multimodal collaborative learning.
@@
    We propose a deep multimodal collaborative learning framework (DMCL) to obtain the visual and texture information simultaneously, and then promote the development of multimodal polyp re-identification.
@@
    A simple but effective multimodal feature fusion strategy is introduced to help the model learn more discriminatve information during training and testing stage.
@@
    Comprehensive experiments on benchmarks demonstrate that our system outperforms state-of-the-art methods of polyp ReID by a clear margin, promoting the development of related methods in clinically assisted colorectal polyp recognition.
@@ 2.1 Hand-crafted based Approaches
Traditional research works (Prosser et al, 2010; Zhao et al, 2013) related to hand-crafted systems for image retrieval task aim to design or learn discriminative representation or pedestrian features. For example, Prosser et al (2010) propose a reformulation of the person re-identification problem as a learning to rank problem. Zhao et al (2013) exploit the pairwise salience distribution relationship between pedestrian images, and solve the person re-identification problem by proposing a salience matching strategy. Unfortunately, these handcrafted feature based approaches always fail to produce competitive results on large-scale datasets. The main reason is that these early works are mostly based on heuristic design, and thus they could not learn optimal discriminative features on current large-scale datasets.
@@ 2.2 Deep Learning based Approaches
Recently, there has been a significant research interest in the design of deep learning based approaches for image or video retrieval (Shao et al, 2021; Lin et al, 2020; Wu et al, 2018). For example, Shao et al (2021) propose temporal context aggregation which incorporates long-range temporal information for content-based video retrieval. Ma et al (2020) explore a method for performing video moment retrieval in a weakly-supervised manner, which learns a sharper attention by pruning out spurious candidate proposals. Lin et al (2020) propose a semantic completion network including the proposal generation module to score all candidate proposals in a single pass. As for the self-supervised learning, Wu et al (2018) present an unsupervised feature learning approach called instance-wise contrastive learning by maximizing distinction between instances via a non-parametric softmax formulation. However, all above approaches divide the learning process into multiple stages, each requiring independent optimization. On the other hand, despite the tremendous successes achieved by deep learning-based approach, they have been left largely unexplored in terms of semantic feature embedding.
To address these challenges, in our case, we employ the visual-text feature with collaborative learning mechanism for colonoscopic polyp ReID task, which not only improves the polyp ReID performance to a certain extent but also effectively maintains the social logical consistency of the extracted feature, ensuring the reliability and usability of the paradigm in various application scenarios. We hope that our method will shed light on some related researches to move forward, especially for multimodality in medical domain.
Figure 3: The overview of our proposed method on polyp re-identification task, which contains two main parts: 1) visual feature backbone and 2) textual feature backbone respectively. Specially, visual feature backbone is composed of CNN network, and textual feature backbone is consisted of Transformer architecture named Bert model. In addition, a multimodal fusion strategy is introduced to mine the mutual benefits between visual feature and texture feature, which can further boost the performance of proposed DMCL framework.
@@ 3.2 Our Proposed DMCL network
In this work, we design a DMCL framework for polyp re-identification in medical scenarios, which includes a visual feature extractor, a textual feature encoder and dynamic multimodal collaborative learning strategy.
@@
Text Encoder. As for the text encoder module, we employ the BERT model (Devlin et al, 2018) to encode the text description of corresponding polyp. In essence, BERT model is a pre-trained language model based on the Transformer architecture, which learns language representations through masked language modeling and next sentence prediction tasks on a large amount of unlabeled text. In this work, given a text description of the polyp with $n$ tokens, BERT will encode it into a $n$ $\times$ 768-dimensions vector matrix for subsequent feature fusion and processing, formulated as $T_{1},T_{2},T_{3},\cdots,T_{n}]^{1\times 768}$. The bidirectional encoder design of BERT enables it to capture contextual information of words, generating more accurate text representations.
Dynamic Multimodal Collaborative Learning. In the general scenarios, there exists an obvious data discrepancy between image and text in our dataset, which inevitably has a negative impact on the model’s performance. To overcome this issue, we creatively introduce a multimodal collaborative learning strategy based on the self-attention mechanism, which is a crucial part of our proposed deep multimodal collaborative learning framework. To be more specific, we obtain the image encoding of $I$ after passing through the image encoder, whose size is then dimensionally reduced from the original 2048-dimension to 768-dimension. Similarly, we can also obtain the text encoding through the text encoder, which is denoted as:
@@
$P=[I,T_{1},T_{2},T_{3},\cdots,T_{n}]$ (2)
where $P$ represents the features generated by the multimodal collaborative learning. Remarkably, multimodal fusion component consists of multiple self-attention layers to accomplish the fusion of visual feature and text feature. After passing through three linear layers, these three vectors are then inputted into the self-attention module. The output from the self-attention module is:
 $\text{ Output }=\text{ Softmax }\left(\frac{\mathbf{QK}^{\mathrm{T}}}{\sqrt{d}}\right)\mathbf{V},d=32$ (3)
where $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ represents the Query, Key, and Value of self-attention mechanism respectively, and $d$ denotes the dimension of the input vector. After multimodal collaborative learning, we can obtain the new updated features $P_{\text{c}}$:
$P_{\text{c}}=\left[I_{c},T_{c1},T_{c2},T_{c3},\ldots,T_{cn}\right]$ (4)
Generally speaking, visual feature is of paramount importance for the task of polyp re-identification. Inspired by this, we choose the feature of $I_{c}$ at the corresponding position of the original image feature, which can be adopted as the testing feature of the corresponding polyp in the testing phase for downstream polyp ReID task. The whole procedure of our collaborative multimodal training system is illustrated in Algorithm 1.
@@
12:  Perform multimodal collaborative learning: $P_{\text{c}}\leftarrow I+T$;
@@
15:Performing inference with model $\theta_{opt}$;
@@ 3.3 Dynamic Network Updating
In essence, many previous works (Wang et al, 2018; Xiang et al, 2024) have been found that performing training with multiple losses has great potential to learn a robust and generalizable ReID model. Inspired by this, we adopt a identification loss and triplet loss as the optimization metric for our DMCL model during the training process. Specially, the triplet-loss function aims to reduce the feature distance between similar polyp images while increasing the feature distance between different polyp images. The identification loss function, on the other hand, is to enhance the model’s ability to recognize and classify polyp categories.
To be more specific, for a single-label $K$ classification task, the identification loss (cross-entropy loss) is written as,
$\mathcal{L}_{ID}=-\frac{1}{\text{ $M_{batch}$ }}\sum_{i=1}^{\text{ $M_{batch}$}}\sum_{j=1}^{K}y_{ij}\log\hat{y_{ij}}$ (5)
@@ 4.1 Datasets and Evaluation Metric
We conduct experiments on several large-scale public datasets, which include Colo-Pair (Chen et al, 2023) for polyp ReID task, and Market-1501 (Zheng et al, 2015), DukeMTMC-reID (Zheng et al, 2017) and CUHK03 dataset (Li et al, 2014) for person ReID tasks.
To be more specific, Colo-Pair (Chen et al, 2023) is the first collection of complete paired colonoscopy sequences, which contains 60 videos from 30 patients, with 62 query video clips and the corresponding polyp clips from the second screening manually annotated as positive retrieval clips.
Market-1501 (Zheng et al, 2015) consists of 32,668 person images of 1,501 identities observed under 6 different camera views. The dataset is split into 12,936 training images of 751 identities and 19,732 testing images of the remaining 750 identities.
DukeMTMC-reID (Zheng et al, 2017) was collected in the winter of Duke University from 8 different cameras, which contains 16,522 images of 702 identities for training, and the remaining images of 702 identities for testing, including 2,228 images as query and 17,661 images as gallery.
CUHK03 (Li et al, 2014) contains 14.096 images which are collected by the Chinese University of Hong Kong with the images taken from only 2 cameras. Specifically, CUHK03 dataset is divided into 7,368 images of 767 identities as the training set and the remaining 5,328 images of 700 identities as the testing set. In this work, we conduct experiments on the labeled bounding boxes (CUHK03 (Labeled)).
In our experiments, we follow the standard evaluation protocol (Zheng et al, 2015) used in the ReID task and adopt mean Average Precision (mAP) and Cumulative Matching Characteristics (CMC) at Rank-1, Rank-5, and Rank-10 for performance evaluation on downstream ReID task.
4.2 Implementation details
In our experiment, ResNet-50 (He et al, 2016) is regarded as the backbone with no bells and whistles during visual feature extraction. Following the training procedure in Xiang et al (2024), we adopt the common methods such as random flipping and random cropping for data augmentation and employ the Adam optimizer with a weight decay co-efficient of $1\times 10^{-5}$ and $1\times 10^{-7}$ for parameter optimization. Besides, we adopt the ID loss and triplet loss functions to train the model for 180 epochs, and the cosine distance is also adopted to calculate the similarity of polyp features in the dataset for the task of polyp re-identification. Please note that the text information is also employed during testing phase. In addition, the batch size $M_{batch}$ for training is set to 64. All the experiments are performed on PyTorch framework with one Nvidia GeForce RTX 2080Ti GPU on a server equipped with an Intel Xeon Gold 6130T CPU.
@@ 4.2 Implementation details
ViSiL (Kordopatis-Zilos et al, 2019) ICCV 24.9 14.5 30.6 51.6
CoCLR (Han et al, 2020) NeurIPS 16.3 6.5 22.6 33.9
TCA (Shao et al, 2021) WCAV 27.8 16.1 35.5 53.2
ViT (Caron et al, 2021) CVPR 20.4 9.7 30.6 43.5
CVRL (Qian et al, 2021) CVPR 23.6 11.3 32.3 53.2
CgSc (Kordopatis-Zilos et al, 2022) IJCV 21.4 8.1 35.5 45.2
@@
Colo-SCRL (Chen et al, 2023) ICME 31.5 22.6 41.9 58.1
VT-ReID (Xiang et al, 2024) ICASSP 37.9 23.4 44.5 60.1